# Guided Query Refinement: Multimodal Hybrid Retrieval with Test-Time Optimization

### v1

@@ Appendix B Model Information
Table 3 depicts the details of the models evaluated in our work. Storage assumes a 16-bit representation and follows official reports (Xu et al., 2025). Jina-Embeddings-v4 supports both text and image document representations, single-vector and multi-vector retrieval, and flexible embedding sizes via a Matryoshka scheme (Kusupati et al., 2022). In its multi-vector textual configuration, the number of vectors per page for Jina varies between pages.
@@ Appendix D Implementation Details
For each model and dataset, we construct an offline index of page-level document representations. For ColPali-based models, multi-page documents are rendered as page images and encoded directly. For text-only embedding models, an ingestion pipeline converts each page to text. We use Docling [^6] https://docling-project.github.io/docling/(Team, 2024) to ingest the images. Docling is an open library providing OCR capabilities combined with document layout analysis, allowing us to recover page content via simple function calls. The resulting text is stored alongside the page images without any chunking, ensuring consistent alignment between visual and textual page representations across the datasets. We run the document converter from Docling v2.342.34 using default parameters.
For Jina-Embeddings-v4 (Günther et al., 2025), a single model accepts both image and text input documents. We thus use it in both a vision configuration and in a text configuration, denoted Jina (Vision) and Jina (Text), both running in a multi-vector setting. Jina (Text) is thus used to test the applicability of GQR where a multi-vector architecture is used as the the GQR complementary encoder.
@@
For RRF we set $\kappa=60$, a common default (Chen et al., 2022; Cormack et al., 2009). For tuning the weighted hybrid baselines and GQR, we follow previous works (Sung et al., 2023; Gangi Reddy et al., 2025; Bruch et al., 2023) and rely on an in-domain development set of queries. We reserve 10% of each subset and tune the hyperparameters, $T$, and $\alpha$ for each, selecting by development set $\mathrm{NDCG}@5$ performance. For the weighted hybrid variants, the weight $\alpha$ is tuned over $\{0.1,0.2,\ldots,0.9\}$ for each subset. For GQR, the learning rates are $\{1\times 10^{-5},\,5\times 10^{-5},\,10^{-4},\,5\times 10^{-4},\,10^{-3},\,5\times 10^{-3}\}$, and we consider $T\in\{10,25,50\}$. The splits are fixed for all experiments. We opted for Adam (Kingma & Ba, 2015) as the optimizer for GQR, based on preliminary experiments comparing different optimizer choices. We use $K=10$ across all retrievers and methods.
Figure 4: Guided Query Refinement (GQR) Architecture. Stage 1: Two retrievers independently encode the query and retrieve top-$K$ documents, forming candidate pool $\mathcal{C}(q)$. Stage 2: complementary retriever produces fixed distribution $p_{1}(\mathcal{C}(q);e^{q}_{1})$ throughout iterations. Primary query embedding $z^{(t)}$ is iteratively refined over $T$ iterations by minimizing KL divergence $\mathcal{L}_{\text{KL}}=\text{KL}(p_{\text{avg}}\|p_{1})$ between consensus distribution $p_{\text{avg}}=\frac{1}{2}(p_{1}(\mathcal{C}(q);z^{(t)})+p_{2}(\mathcal{C}(q);e^{q}_{2}))$ and primary distribution $p_{1}(\mathcal{C}(q);z^{(t)})$. Query representation is updated as $z^{(t+1)}=z^{(t)}-\alpha\nabla_{z^{(t)}}\mathcal{L}_{\text{KL}}$. Final ranking $\mathcal{R}(q)$ uses refined scores from $z^{(T)}$.
Appendix E Additional Results
@@
Table 5 presents the results on ViDoRe 1. It is noticeable that this benchmark suffers from saturation, with many subset scores reaching $90$ or higher (and indeed this was the direct motivation for the release of ViDoRe 2, Macé et al., 2025). Nevertheless, our method does not harm performance, in contrast to other hybrid retrieval methods, as seen in Tables 6 and 7.

### v2

@@ Appendix B Model Information
Table 4 depicts the details of the models evaluated in our work. Storage assumes a 16-bit representation and follows official reports (xu2025llama_nemo). Jina-Embeddings-v4 supports both text and image document representations, single-vector and multi-vector retrieval, and flexible embedding sizes via a Matryoshka scheme (kusupati2022matryoshka). In its multi-vector textual configuration, the number of vectors per page for Jina varies between pages.
@@ Appendix C Additional Definitions
Average Ranking - Tuned (for M=2)
$\mathrm{AvgRank}(d)=\alpha\,\ {rank}_{m_{1}}(d)+(1-\alpha)\,\ {rank}_{m_{2}}(d)$ (9)
RRF - Tuned (for M=2)
$\mathrm{RRF}(d)=\frac{2\alpha}{\,\kappa+\mathrm{rank}_{m_{1}}(d)\,}+\frac{2(1-\alpha)}{\,\kappa+\mathrm{rank}_{m_{2}}(d)\,}$ (10)
Score aggregation - Tuned (for M=2)
$\mathrm{Score}(q,d_{i})=\alpha\,\ \tilde{s}_{m_{1}}(q,d_{i})+(1-\alpha)\,\ \tilde{s}_{m_{2}}(q,d_{i})$ (11)
@@ Appendix D Implementation Details
For each model and dataset, we construct an offline index of page-level document representations. For ColPali-based models, multi-page documents are rendered as page images and encoded directly. For text-only embedding models, for the ViDoRe 1 and ViDoRe 2 datasets an ingestion pipeline converts each page to text. We use Docling [^6] https://docling-project.github.io/docling/(Docling) to ingest the images. Docling is an open library providing OCR capabilities combined with document layout analysis, allowing us to recover page content via simple function calls. The resulting text is stored alongside the page images without any chunking, ensuring consistent alignment between visual and textual page representations across the datasets. We run the document converter from Docling v2.342.34 using default parameters. For ViDoRe 3, we use the official markdown of the benchmark as the text input for the embedding models.
For Jina-Embeddings-v4 (gunther2025jina_v4), a single model accepts both image and text input documents. We thus use it in both a vision configuration and in a text configuration, denoted Jina (Vision) and Jina (Text), both running in a multi-vector setting. Jina (Text) is thus used to test the applicability of GQR where a multi-vector architecture is used as the the GQR complementary encoder.
@@
For RRF we set $\kappa=60$, a common default (chen2022out; 36196). For tuning the weighted hybrid baselines and GQR, we follow previous works (sung-etal-2023-optimizing; 10.1145/3726302.3730160; bruch2023analysis) and rely on an in-domain development set of queries. We reserve 10% of each subset and tune the hyperparameters, $T$, and $\alpha$ for each, selecting by development set $\mathrm{NDCG}@5$ performance. The splits are fixed for all experiments. For the weighted hybrid variants, the weight $\alpha$ is tuned over $\{0.1,0.2,\ldots,0.9\}$ for each subset. For GQR, the learning rates are $\{1\times 10^{-5},\,5\times 10^{-5},\,10^{-4},\,5\times 10^{-4},\,10^{-3},\,5\times 10^{-3}\}$, and we consider $T\in\{10,25,50\}$. We report the selected hyperparameters for all methods in Tables 20-25. We opted for Adam (kingma2015adam) as the optimizer for GQR, based on preliminary experiments comparing different optimizer choices. We use $K=10$ across all retrievers and methods.
Algorithm 1 Guided Query Refinement (GQR)
[…]
6:for $t=0$ to T1T-1 do
11:end for
[…]
14:return $\mathcal{R}(q)$
Appendix E Additional Design Choices for GQR
Candidate pool policy.
Primary and complementary roles
Appendix F Additional Results
Figure 7: Baseline results on ViDoRe 2 across different values of the weight $\alpha$, averaged over 66 model pairs. It shows that the optimal weight for each method varies. Generally the softmax variant emerges as the most stable, while the min-max variant reaches the best performance if tuned correctly. The values for the zero-shot configuration for each baseline in Table 5 were selected based on this plot.
Table 5: Percentage gain in $\mathrm{NDCG}@5$ of hybrid retrieval over the primary retriever on ViDoRe 3. Each cell reports the average gain and standard deviation over 66 retriever pairs (22 multimodal base retrievers and $3$ text retrievers). The Hyperparameters column indicates which hyperparameters were selected for zero-shot evaluation.
Method Avg Std Hyperparameters
[…]
@@ Appendix F Additional Results
Table 7 presents the results on ViDoRe 1. It is noticeable that this benchmark suffers from saturation, with many subset scores reaching $90$ or higher (and indeed this was the direct motivation for the release of ViDoRe 2, mace2025vidore). Nevertheless, our method does not harm performance, in contrast to other hybrid retrieval methods, as seen in Tables 8 and 9.
Storage.
Appendix G Analyzing Per-query Dynamics
As we explain in subsection 2.4, the dynamics in GQR differ from a simple weighted average of scores – the distribution changes are constrained by the geometry of the primary model’s embedding space, and different documents can shift by different magnitudes. To demonstrate this more directly, in Figure 8 we visualize the changes in document ranks for a given query as a function of the number of GQR steps $t$, and as function of the weight $\alpha$ of the Score Aggregation (Min-Max) baseline. As can be seen in the figure, for some queries the dynamics of GQR and score aggregation are relatively similar, whereas for others the effects and dynamics are starkly different. Notably, while score aggregation operates across all candidate documents, the effects of GQR are often focused on a specific subset of documents.
Figure 8: Query-level dynamics of GQR versus score aggregation. The heat maps depict examples of how the document ranks change as a function of either the GQR optimization step $t$ (center left panels), or the weight $\alpha$ of a linear score aggregation baseline (Score Aggregation - Min-Max, center right panels). For both methods we see a gradual progression from the ranks of the primary retriever $m_{1}$ (left) towards those of the complementary retriever $m_{2}$ (right). However, the dynamics differ across queries, as well as across documents within a query. Examples are taken from the ESG subset of ViDoRe 2, for the model pair Colnomic-Embed-Multimodal-7B,Linq-Embed-Mistral (using a GQR learning rate of $10^{-4}$).
@@ Appendix G Analyzing Per-query Dynamics
Table 20: Tuned values for the GQR step size parameter $\alpha$ (learning rate) over ViDoRe 2.
Selected $\alpha$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text $5\times 10^{-4}$ $5\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-4}$
Linq-Embed-Mistral $5\times 10^{-4}$ $5\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-3}$
Qwen3-Embedding-4B $\times 10^{-3} 1×1$$\times 10^{-5}$ $\times 10^{-3} 1×1$$\times 10^{-4}$
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text $\times 10^{-5} 1×1$$\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-4}$
Linq-Embed-Mistral $5\times 10^{-4}$ $\times 10^{-3} 1×1$$\times 10^{-4}$ $1\times 10^{-3}$
Qwen3-Embedding-4B $5\times 10^{-4}$ $1\times 10^{-5}$ $5\times 10^{-5}$ $1\times 10^{-3}$
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text $5\times 10^{-5}$ $\times 10^{-5} 1×1$$\times 10^{-5}$ $1\times 10^{-5}$
Linq-Embed-Mistral $1\times 10^{-4}$ $5\times 10^{-5}$ $\times 10^{-4} 1×1$$\times 10^{-5}$
Qwen3-Embedding-4B $1\times 10^{-4}$ $5\times 10^{-5}$ $\times 10^{-5} 1×1$$\times 10^{-5}$
Table 21: Tuned values for the number of GQR optimization steps parameter $T$ over ViDoRe 2.
Selected $T$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text 10 25 25 25
Linq-Embed-Mistral 25 25 50 10
Qwen3-Embedding-4B 25 10 10 50
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text 10 50 25 25
Linq-Embed-Mistral 25 25 50 25
Qwen3-Embedding-4B 10 50 50 10
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text 25 50 50 10
Linq-Embed-Mistral 10 50 10 10
Qwen3-Embedding-4B 10 10 50 10
Table 22: Tuned values for the weight parameter $\alpha$ of the Average Ranking baseline over ViDoRe 2.
Selected $\alpha$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text 0.8 0.7 0.6 0.8
Linq-Embed-Mistral 0.4 0.5 0.7 0.8
Qwen3-Embedding-4B 0.9 0.9 0.9 0.8
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text 0.9 0.5 0.5 0.8
Linq-Embed-Mistral 0.7 0.3 0.9 0.2
Qwen3-Embedding-4B 0.9 0.9 0.8 0.5
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text 0.8 0.9 0.9 0.9
Linq-Embed-Mistral 0.9 0.5 0.9 0.9
Qwen3-Embedding-4B 0.9 0.9 0.8 0.9
Table 23: Tuned values for the weight parameter $\alpha$ of the RRF baseline over ViDoRe 2.
Selected $\alpha$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text 0.9 0.7 0.6 0.9
Linq-Embed-Mistral 0.5 0.5 0.9 0.7
Qwen3-Embedding-4B 0.9 0.9 0.9 0.6
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text 0.9 0.5 0.5 0.8
Linq-Embed-Mistral 0.7 0.3 0.8 0.6
Qwen3-Embedding-4B 0.9 0.9 0.8 0.8
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text 0.8 0.9 0.9 0.9
Linq-Embed-Mistral 0.9 0.4 0.9 0.9
Qwen3-Embedding-4B 0.9 0.9 0.9 0.9
Table 24: Tuned values for the weight parameter $\alpha$ of the Score Aggregation (Min-Max) baseline over ViDoRe 2.
Selected $\alpha$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text 0.8 0.4 0.7 0.8
Linq-Embed-Mistral 0.6 0.5 0.6 0.8
Qwen3-Embedding-4B 0.7 0.8 0.9 0.7
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text 0.8 0.7 0.6 0.8
Linq-Embed-Mistral 0.7 0.5 0.8 0.6
Qwen3-Embedding-4B 0.8 0.9 0.8 0.6
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text 0.8 0.7 0.8 0.9
Linq-Embed-Mistral 0.7 0.8 0.7 0.9
Qwen3-Embedding-4B 0.8 0.9 0.9 0.9
Table 25: Tuned values for the weight parameter $\alpha$ of the Score Aggregation (Softmax) baseline over ViDoRe 2.
Selected $\alpha$ value
Primary model Complementary model Biomedical Economics ESG ESG Human
Colnomic-Embed-Multimodal-7B Jina-Embeddings-V4 - Text 0.6 0.1 0.5 0.7
Linq-Embed-Mistral 0.1 0.2 0.1 0.1
Qwen3-Embedding-4B 0.1 0.2 0.4 0.1
Jina-Embeddings-V4 Jina-Embeddings-V4 - Text 0.9 0.5 0.6 0.6
Linq-Embed-Mistral 0.2 0.1 0.2 0.1
Qwen3-Embedding-4B 0.6 0.5 0.9 0.1
Llama-Nemoretriever-Colembed-3B-V1 Jina-Embeddings-V4 - Text 0.8 0.8 0.9 0.9
Linq-Embed-Mistral 0.1 0.1 0.1 0.9
Qwen3-Embedding-4B 0.2 0.6 0.8 0.9